a^2+2ab+b^2=25, find a+b
Found 2 solutions by richard1234, fractalier:
Answer by richard1234(7193)

Left-hand side factors to (a+b)^2

(a+b)^2 = 25

a+b = 5 or -5

Answer by fractalier(6550)

As before, we factor
a^2+2ab+b^2=25
and get
(a + b)^2 = 25
Now take the root and get
a + b = 5 and
a + b = -5
Two solutions.